OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent? n/4 = 15/20 A. 2 B. 3 C. 4 D. 5

OpenStudy (danjs):

now this one has a correct answer to it

OpenStudy (danjs):

youc an do it a couple ways, first way , reduce 15 / 20 by dividing both top and bottom by 5

OpenStudy (anonymous):

for 15 divided by 5 is 3, and 20 divided by 5 is 4.

OpenStudy (danjs):

right, so you get \[\frac{ n }{ 4 } = \frac{ 3 }{ 4 }\]

OpenStudy (danjs):

what makes the left look the same as the right

OpenStudy (anonymous):

The 4!

OpenStudy (danjs):

This is the same? \[\frac{ 4 }{ 4 } = \frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

I ment the 3. I thought you ment what looks the same and both have 4.

OpenStudy (danjs):

no, i meant what makes both sides the same, sorry right, n=3

OpenStudy (anonymous):

Thousand thanks to you!!